# Comprehensive benchmarking of large language models for RNA secondary structure prediction

### v1

@@ Results
Figure 3: |||| Inter-family structure prediction based on RNA LLM. Performance evaluated on the 9 RNA families of the ArchiveII dataset. Each boxplot represents the $F_{1}$ performance of all methods for a given family in the test set. The median performance of LinearPartition-C is shown, for each family, as a horizontal blue line. Below RNA-LLM performances, sequence length distribution for each dataset is shown in blue and distribution of minimum test-train structural distance [38] is shown in orange. a, tRNA family. A Wilcoxon test for paired samples with Bonferroni correction indicates that all differences are statistically significant ($P<0.0001$, $N=557$), except in the case of RiNALMo with ERNIE-RNA, and RNABERT with one-hot. b, 5s family. All differences are statistically significant ($P<0.0001$, $N=1283$). c, tmRNA family. All differences are statistically significant ($P<0.0001$, $N=462$). d, RNaseP family. Only ERNIE-RNA with RiNALMo, RNA-MSM with RNAErnie, and one-hot with RNABERT are not statistically significant. e, srp family. Differences for the top 4 methods are statistically significant ($P<0.0001$, $N=918$). f, grp1 family. Differences for the top 2 methods are statistically significant ($P<0.0001$, $N=74$). g, 23s family. Top 2 methods are significantly better than the rest of all the methods ($P<0.01$, $N=15$). h, 16s family. RiNALMo and ERNIE-RNA are statistically different and are also statistically different from the rest of the methods ($P<0.0001$, $N=66$). i, telomerase family. ERNIE-RNA and one-hot are statistically different from RNABERT, RiNALMo, RNA-MSM and RNAErnie. Details of statistical analysis in Supplementary Fig. S2.
@@
In the case of the tRNA family (Fig. 3a), the one with the shortest sequence length as shown in the panel below, most methods achieve performance above $F_{1}=0.40$. RNABERT and one-hot achieved both the same performance here, close to $F_{1}=0.46$. The four best methods for this family are RiNALMo ($F_{1}=0.86$), ERNIE-RNA ($F_{1}=0.84$), RNAErnie ($F_{1}=0.81$) and RNA-FM ($F_{1}=0.79$). In this case, which can be considered the easiest one for a cross-family validation due to the small length of sequences and test/train structural distance below 0.5, four of the RNA-LLM outperform the classical LinearPartition-C, which achieved $F_{1}=0.75$. For the 5s family (Fig. 3b), the results for all RNA-LLM are very similar to the previous family, again due to the length of the sequences (the second mean shortest family) and low structural distance between this family and all the remaining families in the training set. In this case, one-hot representation, RNABERT and RNA-MSM are below $F_{1}=0.26$, while RNAErnie and RNA-FM achieve performances around $F_{1}=0.55$, and again ERNIE-RNA and RiNALMo achieved both the best results, with $F_{1}=0.84$ and $F_{1}=0.82$, respectively, above the classical method ($F_{1}=0.78$). For the tmRNA (Fig. 3c) and RNaseP (Fig. 3d) families, the conclusions are very similar as before. Although ERNIE-RNA and RiNALMo have performance 10% lower than with the previous families, in these cases achieved even higher performance than the classical method. In the case of the tmRNA family (Fig. 3c), the classical method has $F_{1}=0.38$, while ERNIE-RNA and RiNALMo have almost twice the performance with $F_{1}=0.73$. In the case of the RNaseP (Fig. 3d) family, those two RNA-LLM methods are more than 10% better than the classical one. Here, it is important to note that these two families are very similar to each other in average sequence length, around 300 nt, but 3 times longer than tRNA and 5s families.
@@
In the case of the srp family (Fig. 3e), there is an important change in the length distribution, which becomes bimodal, and in the structural distances distribution, which clearly exceeds the 0.5 mark. Here, the classical method, ERNIE-RNA and RiNALMo achieve extremely close results while all other RNA-LLM are far below them. ERNIE-RNA has median $F_{1}=0.67$ while the classical method achieved $F_{1}=0.68$. Notably, in the case of the grp1 family (Fig. 3f), which has larger average length size than srp, RiNALMo is the only LLM that slightly outperforms the classical method ($F_{1}=0.64$ versus $F_{1}=0.61$). In the case of the 23s and 16s families (Figs. 3g and 3h, respectively), it can be seen that the one-hot representation, RNABERT, RNA-FM, RNA-MSM and RNAErnie embeddings are all below $F_{1}=0.25$, a very low performance. However, ERNIE-RNA and RiNALMo achieve almost twice that performance with median values near to $F_{1}=0.60$, and the classical method achieves slightly better results ($F_{1}=0.69$ for 23s and $F_{1}=0.66$ for 16s). Notably, this last family has a wide variety of sequence lengths and structural distances. Finally when the telomerase family (Fig. 3i) is used in the test set, all RNA-LLM achieve extremely poor results below $F_{1}=0.15$. This family has the longest sequences. Here the classical method has a large advantage over all RNA-LLM, although it is low compared to the other families.
@@ Discussion
We benchmarked 6 RNA-LLM for their ability to predict RNA secondary structure structure based only on the sequence, in 4 datasets of increasing complexity. First we performed a visual comparison of the LLM embeddings with a UMAP projection, analyzing the distribution and separation of the RNA families. We found that RiNALMo and ERNIE-RNA were the models that could better represent and separate the RNA families in the projection without almost overlap. Notably, as Table 1 shows, those RNA-LLM are the biggest ones, pre-trained with the largest and most varied datasets.
@@
We tested each RNA-LLM with exactly the same experimental setup and deep neural network architecture for RNA secondary structure prediction in several benchmark datasets, increasing the difficulty from simple random partitions to very low homology between training and testing partitions. Overall, RiNALMo and ERNIE-RNA maintained an acceptable level of generalization capability even in the hardest test, providing equivalent performance to the classical folding method. Remarkably, in that hardest scenario the trivial one-hot representation achieved a performance similar to the best performing RNA-LLM. Thanks to the transfer learning paradigm, one would have expected that LLM would help to generalize even with a small training set and prediction model, because all the relevant information was learnt in the embeddings. That was not the case and it was probably an indication that there is still important information that cannot be fully captured even by the LLM pre-training. The results achieved might be explained by the fact that the structures of sequences in the testing sets were really very different from those seen during pre-training, or they were a small minority, thus the generalization capability of the prediction model could not be really benefited from using a LLM representation.
@@
For the most difficult secondary structure prediction task, we removed from training all the sequences of one specific testing family. In this setup, most of the RNA-LLM achieved a high performance for the two shortest families. However, as the average length of the sequence to be predicted increased, the performance of most RNA-LLM lowered below $F_{1}=0.50$. Overall, in 4 out of 9 cases RiNALMo and ERNIE-RNA outperformed the classical method. In two cases no differences were observed, and in the last three families the classical method was the best. In the case of the telomerase RNA family, the most different from the other ones, with very few samples in all datasets and the largest average sequence length, the results were all poor, up to three times below the classical method. In any case, in this RNA family the classical method also obtained a performance without practical usefulness, barely reaching $F_{1}=0.50$.
@@
Acknowledgments
This work was supported by Argentine National Scientific and Technical Research Council (CONICET), ANPCyT (PICT 2022–0086) and UNL (CAI+D 2020 115). We also acknowledge the support of NVIDIA Corporation for the donation of GPU used for this research.
Author contributions
GS and DHM conceived the idea, designed and initiated the study. DHM coordinated collaborations and communication between team members. LZ, LB and DHM developed the experimental design and source code. GS wrote the first manuscript. GS and DHM edited the manuscript with contributions from all coauthors. DHM and MG created figures. LZ, LB and DHM worked on data visualization. LZ and LB implemented the deep learning model and built the repositories. LB and LZ curated and managed the datasets. LDP, GS and DHM performed statistical analyses. All the authors perform computational evaluation of models. All authors reviewed and approved the final manuscript.

### v2

@@ Results
Figure 3: Inter-family structure prediction based on RNA LLM. Performance evaluated on the 9 RNA families of the ArchiveII dataset. Each boxplot represents the $F_{1}$ performance of all methods for a given family in the test set. The thermodynamic prediction method (LinearPartition-V, dashed black) and the DL-based method (sincFold, solid black line) are added as baselines in all plots. Below RNA-LLM performances, sequence length distribution for each dataset is shown in blue and distribution of minimum test-train structural distance [46] is shown in orange. a, tRNA family. A Wilcoxon test for paired samples with Bonferroni correction indicates that all differences are statistically significant ($P<0.0001$, $N=557$), except in the case of RiNALMo with ERNIE-RNA, and RNABERT with one-hot. b, 5s family. All differences are statistically significant ($P<0.0001$, $N=1283$). c, tmRNA family. All differences are statistically significant ($P<0.0001$, $N=462$). d, RNaseP family. Only ERNIE-RNA with RiNALMo, RNA-MSM with RNAErnie, and one-hot with RNABERT are not statistically significant. e, srp family. Differences for the top 4 methods are statistically significant ($P<0.0001$, $N=918$). f, grp1 family. Differences for the top 2 methods are statistically significant ($P<0.0001$, $N=74$). g, 23s family. Top 2 methods are significantly better than the rest of all the methods ($P<0.01$, $N=15$). h, 16s family. RiNALMo and ERNIE-RNA are statistically different and are also statistically different from the rest of the methods ($P<0.0001$, $N=66$). i, telomerase family. ERNIE-RNA and one-hot are statistically different from RNABERT, RiNALMo, RNA-MSM and RNAErnie. Details of statistical analysis in Supplementary Fig. S2.
For assessing inter-family performance, a family-fold cross validation in the ArchiveII dataset was performed. That is, one family was left out for testing per cross-validation fold, and the rest of the families were used for training. This eliminates most of the homology to the training set, providing a hard measure of performance and, thus, allowing estimating future performance on novel RNAs that do not belong to any known family. These results are shown in Figure 3, along with the distributions of sequence lengths and minimum structural distances [46] between each test sequence and all the sequences in its corresponding training fold. The best thermodynamic and DL-based predictors are shown as reference for comparison in the plots (detailed results in Supplementary Table 2).
@@
In the case of the tRNA family (Fig. 3a), the one with the shortest sequence length as shown in the panel below, most methods achieve performance above $F_{1}=0.40$. RNABERT and one-hot achieved both the same performance here, close to $F_{1}=0.46$. The four best methods for this family are RiNALMo ($F_{1}=0.86$), ERNIE-RNA ($F_{1}=0.84$), RNAErnie ($F_{1}=0.81$) and RNA-FM ($F_{1}=0.79$). In this case, which can be considered the easiest one for a cross-family validation due to the small length of sequences and test/train structural distance below 0.5, four of the RNA-LLM outperform both the classical method, which achieved $F_{1}=0.70$, and the DL-based model, which achieved $F_{1}=0.71$. For the 5s family (Fig. 3b), the results for all RNA-LLM are very similar to the previous family, again due to the length of the sequences (the second mean shortest family) and low structural distance between this family and all the remaining families in the training set. In this case, one-hot representation, RNABERT and RNA-MSM are below $F_{1}=0.26$, while RNAErnie and RNA-FM achieve performances around $F_{1}=0.55$, and again ERNIE-RNA and RiNALMo achieved both the best results, with $F_{1}=0.84$ and $F_{1}=0.82$, respectively, above the classical method ($F_{1}=0.73$). The DL-based model achieved $F_{1}=0.46$ here, below RNAErnie and RNA-FM. For the tmRNA (Fig. 3c) and RNaseP (Fig. 3d) families, the conclusions are very similar as before. Although ERNIE-RNA and RiNALMo have performance 10% lower than with the previous families, in these cases achieved even higher performance than both the classical and the DL-based method. In the case of the tmRNA family (Fig. 3c), the classical method has $F_{1}=0.42$ and the DL-based has $F_{1}=0.34$, while ERNIE-RNA and RiNALMo have almost twice the performance with $F_{1}=0.73$. In the case of the RNaseP (Fig. 3d) family, those two RNA-LLM methods are more than 10% better than the classical and the DL-based method. Here, it is important to note that these two families are very similar to each other in average sequence length, around 300 nt, but 3 times longer than tRNA and 5s families.
@@
In the case of the srp family (Fig. 3e), there is an important change in the length distribution, which becomes bimodal, and in the structural distances distribution, which clearly exceeds the 0.5 mark. Here, the classical method, ERNIE-RNA and RiNALMo achieve extremely close results while all other RNA-LLM and the DL-based method are far below them. ERNIE-RNA has median $F_{1}=0.67$ while the classical method achieved $F_{1}=0.66$. Notably, in the case of the grp1 family (Fig. 3f), which has larger average length size than srp, RiNALMo is the only LLM that clearly outperforms all methods, even the classical one ($F_{1}=0.64$ versus $F_{1}=0.57$). In the second position, ERNIE-RNA achieves the same performance as the DL-based method. In the case of the 23s and 16s families (Figs. 3g and 3h, respectively), it can be seen that the one-hot representation, RNABERT, RNA-FM, RNA-MSM and RNAErnie embeddings are all below $F_{1}=0.25$, a very low performance. However, ERNIE-RNA and RiNALMo achieve almost twice that performance with median values near to $F_{1}=0.60$, and the classical method achieves better results ($F_{1}=0.74$ for 23s and $F_{1}=0.58$ for 16s). The DL-based method presents intermediate results: $F_{1}=0.40$ for 23s and $F_{1}=0.41$ for 16s.
Notably, this last family has a wide variety of sequence lengths and structural distances. Finally when the telomerase family (Fig. 3i) is used in the test set, all RNA-LLM and the DL-based method achieve extremely poor results below $F_{1}=0.15$. This family has the longest sequences. Here the classical method has a large advantage over all RNA-LLM ($F_{1}=0.45$), although it is low compared to the other families.
For these dataset partitions we have also calculated the WL metric, showing performance trends similar to those analyzed for $F_{1}$ score (see details in Supplementary Table 2).
@@ Discussion
We benchmarked 6 RNA-LLM for their ability to predict RNA secondary structure based only on the sequence, in 4 datasets of increasing complexity. First we performed a visual comparison of the LLM embeddings with a UMAP projection, analyzing the distribution and separation of the RNA families. We found that RiNALMo and ERNIE-RNA were the models that could better represent and separate the RNA families in the projection without almost overlap. Notably, as Table 1 shows, those RNA-LLM are the biggest ones, pre-trained with the largest and most varied datasets.
@@
We tested each RNA-LLM implementation with exactly the same experimental setup and deep neural network architecture for RNA secondary structure prediction in several benchmark datasets, increasing the difficulty from simple random partitions to very low homology between training and testing partitions. Overall, RiNALMo and ERNIE-RNA maintained an acceptable level of generalization capability even in the hardest test, providing equivalent performance to the classical folding method. Remarkably, in that hardest scenario the trivial one-hot representation achieved a performance similar to the best performing RNA-LLM. Thanks to the transfer learning paradigm, one would have expected that LLM would help to generalize even with a small training set and prediction model, because all the relevant information was learnt in the embeddings. That was not the case and it was probably an indication that there is still important information that cannot be fully captured even by the LLM pre-training. The results achieved might be explained by the fact that the structures of sequences in the testing sets were really very different from those seen during pre-training, or they were a small minority, thus the generalization capability of the prediction model could not be really benefited from using a LLM representation.
Precisely regarding the information that may have been incorporated in the pre-training of the RNA-LLMs, it is interesting to compare them with methods that use homology information for predicting RNA secondary structures [5, 21, 73, 51, 22, 69]. In the case of the dataset closest to a real-world application, PDB-RNA (Figure 2d), Singh et al. [55] reviewed the most relevant methods that use homology information with the following performance: CentroidAlifold [21] $F_{1}=0.69$, Turbofold II [65] $F_{1}=0.64$, RNAalifold [5] $F_{1}=0.66$ and CentroidFold [49] $F_{1}=0.60$. These performances are extremely close to the baseline performance of the pure thermodynamic method ($F_{1}=0.67$, Figure 2d), and those are actually the highest performance values for this dataset. Instead, here all RNA-LLM show very poor results, with median performances below $F_{1}=0.40$. Thus, there is still a large performance gap between RNA LLM based and classical methods, independently of explicit homology incorporation. Remarkably, the best DL-based method, without using any homology nor alignment data, has achieved an intermediate performance within the gap mentioned ($F_{1}=0.58$), even very close to some of the homology-based methods. It is also interesting to compare RNA-LLM trained only with RNA sequences with approaches that incorporate other information in pre-training. Multimodal LLM are interesting since they propose new ways of gaining insights in related data distributions. For example, Evo [41] was trained with DNA, considering coding and non-coding sequences. However, since this is fundamentally a generative model trained for sequence design with large contexts, the internal representations obtained for RNA sequences from our datasets were not successful to predict secondary structures. Other methods use abundant DNA sequences in a BERT-like architecture such as DNABERT2 [79]. In this case, the performance of the DNABERT2 model in the datasets here evaluated was just slightly better than RNABERT. From these results, it can be concluded that LLM are still unable to correctly represent the additional information incorporated in the pre-training stage, or if they were able to capture it, it is not trivial to extract this extra information with the models in the downstream tasks.
@@
For the most difficult secondary structure prediction task, we removed from training all the sequences of one specific testing family. In this setup, most of the RNA-LLM achieved a high performance for the two shortest families. However, as the average length of the sequence to be predicted increased, the performance of most RNA-LLM lowered below $F_{1}=0.50$. Overall, in 4 out of 9 cases RiNALMo and ERNIE-RNA outperformed the classical method. In two cases no differences were observed, and in the last three families the classical method was the best. In the case of the telomerase RNA family, the most different from the other ones, with very few samples in all datasets and the largest average sequence length, the results were all poor, up to three times below the classical method. In any case, in this RNA family the classical method also obtained a performance without practical usefulness, barely reaching $F_{1}=0.50$. Remarkably, in this hard prediction task, RiNALMo and ERNIE-RNA always outperformed the DL-based prediction method.